# Shop-R1: Rewarding LLMs to Simulate Human Behavior in Online Shopping via Reinforcement Learning

### v1

@@ Abstract
Large Language Models (LLMs) have recently demonstrated strong potential in generating ‘believable human-like’ behavior in web environments. Prior work has explored augmenting training data with LLM-synthesized rationales and applying supervised fine-tuning (SFT) to enhance reasoning ability, which in turn can improve downstream action prediction. However, the performance of such approaches remains inherently bounded by the reasoning capabilities of the model used to generate the rationales. In this paper, we introduce Shop-R1, a novel reinforcement learning (RL) framework aimed at enhancing the reasoning ability of LLMs for simulation of real human behavior in online shopping environments Specifically, Shop-R1 decomposes the human behavior simulation task into two stages: rationale generation and action prediction, each guided by distinct reward signals. For rationale generation, we leverage internal model signals (e.g., logit distributions) to guide the reasoning process in a self-supervised manner. For action prediction, we propose a hierarchical reward structure with difficulty-aware scaling to prevent reward hacking and enable fine-grained reward assignment. This design evaluates both high-level action types and the correctness of fine-grained sub-action details (attributes and values), rewarding outputs proportionally to their difficulty. Experimental results show that our method achieves a relative improvement of over 65% compared to the baseline.  [^1] The code and model checkpoints will be released upon paper acceptance.
Figure 1: Overview of the proposed reinforcement learning framework, Shop-R1, designed to simulate real human behaviors in web-based shopping environments. Given an action history with corresponding website observations, the model predicts the next action and its rationale based on the history and the latest website observation. The generated responses are evaluated from four perspectives: format correctness, self-certainty of the rationale, action type accuracy, and sub-action (attribute and value) accuracy.
@@ 1 Introduction
Since RL offers a flexible and effective training paradigm, particularly suited for settings with sparse and delayed feedback, and allows for fine-grained control over behavioral outputs [22, 23, 24, 25, 26], we utilize RL for the simulation of human shopping behavior, compared to previous work that focuses primarily on task completion [27, 28], In this work, we propose Shop-R1, a novel RL framework designed to enhance LLMs for simulation of human online shopping behaviors. As shown in Fig. 1, Shop-R1 decomposes the human behavior simulation task into two stages: (1) rationale generation and (2) action prediction, with tailored reward signals for each component. For the reward design, we begin by introducing a binary format reward that encourages the model to produce responses in a parse-friendly structure, thereby facilitating reliable downstream evaluation and reward computation. Specifically, the model receives a non-zero reward only when its output conforms to the expected format; otherwise, it is penalized with zero reward. For rationale generation, acquiring ground-truth rationales is inherently difficult. Although efforts like OPeRA [29] attempt to collect self-reported rationales from real users, such annotations may omit implicit or unconscious decision factors. To address this, we incorporate a self-certainty reward [30, 31], quantified via the average Kullback–Leibler (KL) divergence [32] between the model’s output distribution and a uniform distribution. This signal captures the model’s confidence in its generated rationales, providing a supervision-free alternative to ground-truth rationales. For action prediction, we go beyond binary reward signals by introducing a hierarchical reward scheme that accounts for both action type and sub-action correctness. This design allows the agent to receive partial credit for plausible but imperfect behaviors, promoting smoother and more robust learning. Furthermore, to mitigate reward hacking and reflect the varying difficulty of different actions, we apply a difficulty-aware reward scaling strategy that adjusts the reward magnitude based on action complexity. Our main contributions are summarized as follows:
  • To the best of our knowledge, we are the first to introduce RL into a simulation-oriented human behavior modeling task. We reformulate human online shopping behavior simulation as a two-stage prediction problem, comprising rationale generation and action prediction, and design distinct RL objectives for each.
@@ 2 Related Work
LLM for human behavior simulation. Large Language Models (LLMs) have emerged as powerful tools to simulate human behaviors in diverse real-world settings. Recent advances have led to the development of agent systems capable of generating plausible user actions based on static personas and interaction histories, enabling the modeling of behavior in contexts such as social science [33, 34], recommender systems [35], and user experience research [36]. These systems typically condition on user profiles (e.g., preferences, demographics) and session histories (e.g., clickstreams, task sequences) to predict the next likely user action, allowing for personalized and context-aware simulations. Beyond behavior prediction, recent efforts have enriched these simulations by incorporating explicit reasoning processes. Methods like ReAct [1] and reflection-based models [37, 38] prompt LLMs to produce intermediate thought traces before action generation, enhancing interpretability and decision quality. Systems such as WebAgent [39] and UX-Agent [36] further decompose tasks into sub-goals using dedicated reasoning models, yielding improved control in complex environments like web interfaces. A parallel line of research explores agent-based LLM frameworks that simulate multi-agent interactions in dynamic environments [40, 41, 42]. These systems often adopt modular roles (e.g., planners, executors) and collaborative reasoning [43, 44], offering insights into emergent social behaviors and teamwork dynamics. Despite recent advances, there remains a significant gap in exploring how RL can be leveraged to further enhance the simulation of human behavior using LLMs, particularly in the context of web-based shopping environments.
@@
Reward design for RL. Reward design plays a central role in the effectiveness and generalization of RL algorithms, particularly in the context of aligning LLMs with desired behaviors. The prominent paradigm is Reinforcement Learning from Human Feedback (RLHF), which has been widely adopted to fine-tune LLMs using reward models trained on human preference data [45]. While RLHF has demonstrated strong alignment capabilities, it is often bottlenecked by the high cost and limited scalability of collecting reliable human annotations [46]. Moreover, reward models themselves can introduce alignment biases and inaccuracies, especially when trained on limited or noisy preference comparisons [47]. To alleviate these limitations, Direct Preference Optimization (DPO) [23] proposes a more efficient alternative that directly optimizes model parameters against human preference signals without an explicit reward model. Though computationally lighter, DPO and its variants still depend on the availability and quality of human-generated or approximated preference data, which can be inconsistent across tasks and domains. A complementary direction has emerged through Reinforcement Learning with Verifiable Rewards (RLVR), particularly suited for domains with deterministic correctness criteria such as code generation and mathematical reasoning [48, 49]. RLVR frameworks employ rule-based verifiers to automatically compute reward signals based on strict correctness (e.g., exact string matching or functional equivalence) bypassing the need for human feedback. This shift toward automated objective reward functions has enabled the training of highly capable models such as DeepSeek-R1 [48] and inspired new policy optimization methods such as GRPO [50] and its recent extensions [51, 52]. Despite these advances, reward design remains a fundamental challenge in RL for human behaviors. RLHF offers flexibility for modeling subjective tasks, but often suffers from scalability and reliability issues [53, 54, 55, 56]. In contrast, RLVR provides high precision by relying on clearly defined evaluation criteria, but is limited to tasks where such criteria exist [49, 57, 58]. To address the unique challenges of simulating human online shopping behavior, we propose a hybrid reward framework specifically tailored to this domain.
@@ 3 Methodology
In this section, we first formulate the problem of human behavior simulation in the context of web-based shopping. We then present the design of our proposed RL framework, Shop-R1, tailored specifically for simulating human behavior in this setting.
@@
Problem statement. In the context of web shopping, a user session is composed of a sequence of multi-step actions $a_{1\dots t\dots N}$, typically initiated with a search query and concluded by either a product purchase or a termination action (e.g., closing the browser). Following the setup of [20], the action space comprises three primary action types: ‘type_and_submit’, ‘click’, and ‘terminate’. More details about the action space can be found in App. A. Each action $a_{t}$ is paired with a corresponding rationale $r_{t}$, which captures the user’s underlying motivation or rationale at that time step. The model also receives contextual information, i.e., the observation space, representing the current state of the web environment. This context is encoded as a simplified HTML structure, as introduced in Lu et al. [13], which preserves essential layout and content elements while discarding non-informative components such as scripts and styles. The task of human online shopping behavior simulation is formally defined as learning a function $f$ that predicts the next rationale and action, given the cumulative context and action history:
$f(c_{1\dots t},a_{1\dots t-1},r_{1\dots t-1})=r_{t},a_{t},$ (1)
@@
where $f$ denotes the model trained to simulate user behavior by generating the next-step rationale $r_{t}$ and action $a_{t}$ conditioned on prior context $c_{1\dots t}$, past actions $a_{1\dots t-1}$, and prior rationales $r_{1\dots t-1}$. These rationales are generated using LLMs and serve as supervision signals during the supervised fine-tuning stage for a cold start. Need to note that no generated rationales are used during the subsequent RL stage.
@@
Cold start with SFT. Following the approach of Guo et al. [48], we initialize the behavior simulation model $f$ through supervised fine-tuning (SFT) on annotated trajectories, where each rationale is generated by Claude 3.5 Sonnet [21] via Amazon Bedrock, without leveraging any user profile information. This SFT phase acts as a cold start for subsequent RL, grounding the model in realistic rationale and action patterns. During this phase, the model is trained to jointly generate rationales and corresponding actions. The training objective is to maximize the likelihood of the ground truth rationale-action pairs, conditioned on the the input query $q_{t}=c_{1\dots t},a_{1\dots t-1},r_{1\dots t-1}$:
@@
Shop-R1. To better guide policy optimization in the human behavior simulation setting, we decompose each step into two sub-tasks: rationale generation and action prediction. Each sub-task is assigned a tailored reward to improve alignment and interpretability. To ensure the ease and correctness of parsing predicted rationales and actions from model outputs, we introduce a binary format reward, which encourages the model to produce responses in a structured JSON format. This format adheres to a dictionary schema with two keys: rationale and action. For rationale generation, we employ a self-certainty score [30, 31], which quantifies the model’s confidence in its generated rationale. Specifically, we compute the KL divergence between the model’s predictive distribution over the vocabulary and a uniform distribution, averaged over the entire output sequence:
@@
where $N$ is the number of tokens in the generated rationale $r_{t}$, $p_{ij}$ is the predicted probability of token $i$ at position $j$, and $U_{i}=\frac{1}{|V|}$ is the uniform distribution over the vocabulary $V$. Higher values of $s(\cdot)$ indicate greater certainty and consistency in the model’s reasoning. For action prediction, we replace the brittle binary signal with a hierarchical reward scheme that credits both the coarse-grained action type and its fine-grained sub-actions to stabilize training and discourage degenerate reward hacked policies. This hierarchical scheme densifies the reward landscape: it expands the set of profitable trajectories, lifts the agent out of the ‘no-reward’ plateau that typically stalls policy search, and makes reward hacking uneconomical. Concretely, every action, easy or hard, earns the same coarse-level reward once its high-level type is correct; only the more complex actions can unlock additional gains through their long-text sub-components. As a result, naively spamming the trivial ‘terminate’ action no longer yields a competitive payoff, while executing the full (‘click’, ‘type_and_submit’) sequence becomes the most lucrative strategy. Concretely, a ‘click’ action containing a sub-action, specifying the button name to be clicked; partial rewards are granted for the correctly predicted components. Likewise, ‘type_and_submit’ contains sub-action, providing the intended textual content. In contrast, ‘terminate’ has no sub-actions and is scored only at the action-type level. Prediction accuracy is measured with task-specific metrics: discrete action types use an exact-match criterion, whereas free-form sub-actions are evaluated with ROUGE-L. A text-based sub-action, such as a button label or search query, earns a soft reward proportional to its ROUGE-L similarity to the ground truth, but only when that similarity exceeds a preset threshold (e.g., 0.75). Because long-text sub-actions are substantially harder, where modern webpages can expose thousands of candidate elements, we introduce a difficulty-aware reward scaling (DARS) factor that amplifies rewards for correctly predicting these components. This prevents reward hacking behaviors in which the agent repeatedly selects the trivial ‘terminate’ action to secure easy points. The proposed hierarchical reward scheme is summarized in Tab. 1. Bringing these components together, the objective of Shop-R1is to maximize the combined reward signal derived from multiple sources, while regularizing with a KL divergence to a reference policy:
where $\pi_{\text{ref}}$ denotes a fixed reference policy, $v(a_{t})$ denotes the reward for action prediction and $\alpha$ and $\beta$ are hyperparameters that control the strength of the corresponding regularization terms.
@@
[…]

### v2

@@ Abstract
Large Language Models (LLMs) have recently demonstrated strong potential in generating ‘believable human-like’ behavior in web environments. Prior work has explored augmenting training data with LLM-synthesized rationales and applying supervised fine-tuning (SFT) to enhance reasoning ability, which in turn can improve downstream action prediction. However, the performance of such approaches remains inherently bounded by the reasoning capabilities of the model used to generate the rationales. In this paper, we introduce Shop-R1, a novel reinforcement learning (RL) framework aimed at enhancing the reasoning ability of LLMs for simulation of real human behavior in online shopping environments. Specifically, Shop-R1 decomposes the human behavior simulation task into two stages: rationale generation and action prediction, each guided by distinct reward signals. For rationale generation, we leverage internal model signals (e.g., logit distributions) to guide the reasoning process in a self-supervised manner. For action prediction, we propose a hierarchical reward structure with difficulty-aware scaling to prevent reward hacking and enable fine-grained reward assignment. This design evaluates both high-level action types and the correctness of fine-grained sub-action details (attributes and values), rewarding outputs proportionally to their difficulty. Experimental results show that our method achieves a relative improvement of over 65% compared to the baseline. The project page is available at https://damon-demon.github.io/shop-r1.html.
Figure 1: Overview of the proposed reinforcement learning framework, Shop-R1, designed to simulate real human behaviors in web-based shopping environments. Given an action history $a_{1\dots t-1}$ with corresponding website observations $c_{1\dots t-1}$, the model predicts the next action $a_{t}$ and its rationale $r_{t}$ based on the history and the latest website observation $c_{t}$. The generated responses are evaluated from four perspectives: format correctness, self-certainty of the rationale, action type accuracy, and sub-action (attribute and value) accuracy.
@@ 1 Introduction
Since RL offers a flexible and effective training paradigm, particularly suited for settings with sparse and delayed feedback, and allows for fine-grained control over behavioral outputs (Kaufmann et al., 2024; Rafailov et al., 2023; Mu et al., 2024; Bai et al., 2022; Glaese et al., 2022), we utilize RL for simulating human shopping behavior, in contrast to prior work that focuses primarily on task completion (Lin et al., 2022; Zhou et al., 2023; Dong et al., 2025), where the user goal is clearly specified. In this work, we propose Shop-R1, a novel RL framework with hierarchical reward scheme designed to enhance LLMs for simulation of human online shopping behaviors. As shown in Fig.˜1, Shop-R1 decomposes the human behavior simulation task into two stages: (1) rationale generation and (2) action prediction, with tailored reward signals for each component, since reasoning before action prediction has been shown to improve performance (Zhang et al., 2024a). For the reward design, we begin by introducing a binary format reward that encourages the model to produce responses in a parse-friendly structure, thereby facilitating reliable downstream evaluation and reward computation. Specifically, the model receives a non-zero reward only when its output conforms to the expected format; otherwise, it is penalized with zero reward. A more detailed discussion on the choice between zero and negative rewards is provided in Appx. A. For rationale generation, acquiring ground-truth rationales is inherently difficult. Although efforts like OPeRA (Wang et al., 2025c) attempt to collect self-reported rationales from real users, such annotations may omit implicit or unconscious decision factors. To address this, we incorporate a self-certainty reward (Kang et al., 2025; Zhao et al., 2025), quantified via the average Kullback–Leibler (KL) divergence (Csiszár, 1975) between the model’s output distribution and a uniform distribution. This signal captures the model’s confidence in its generated rationales, providing a supervision-free alternative to ground-truth rationales that contributes to training stability. For action prediction, we go beyond binary reward signals by introducing a hierarchical reward scheme that accounts for both action type and sub-action correctness. This design allows the agent to receive partial credit for plausible but imperfect behaviors, promoting smoother and more robust learning. Furthermore, to mitigate reward hacking and reflect the varying difficulty of different actions, we apply a difficulty-aware reward scaling strategy that adjusts the reward magnitude based on action complexity. Our main contributions are summarized as follows:
  • To the best of our knowledge, we are the first to introduce RL into a simulation-oriented human behavior modeling task in web-shopping environments. We reformulate human online shopping behavior simulation as a two-stage prediction problem, comprising rationale generation and action prediction, and design distinct RL objectives for each.
@@ 2 Related Work
LLM for human behavior simulation. Large Language Models (LLMs) have emerged as powerful tools to simulate human behaviors in diverse real-world settings. Recent advances have led to the development of agent systems capable of generating plausible user actions based on static personas and interaction histories, enabling the modeling of behavior in contexts such as social science (Park et al., 2023b; 2024), recommender systems (Wang et al., 2023), e-commerce (Wang et al., 2025d), and user experience research (Lu et al., 2025d). These systems typically condition on user profiles (e.g., preferences, demographics) and session histories (e.g., clickstreams, task sequences) to predict the next likely user action, allowing for personalized and context-aware simulations. Beyond behavior prediction, recent efforts have enriched these simulations by incorporating explicit reasoning processes. Methods like ReAct (Yao et al., 2023) and reflection-based models (Shinn et al., 2023; Park et al., 2023a) prompt LLMs to produce intermediate thought traces before action generation, enhancing interpretability and decision quality. Systems such as WebAgent (Gur et al., 2023) and UX-Agent (Lu et al., 2025d) further decompose tasks into sub-goals using dedicated reasoning models, yielding improved control in complex environments like web interfaces. A parallel line of research explores agent-based LLM frameworks that simulate multi-agent interactions in dynamic environments (Ma et al., 2024; Wang et al., 2025b; OpenAI, 2025). These systems often adopt modular roles (e.g., planners, executors) and collaborative reasoning (Qian et al., 2024; Luo et al., 2024), offering insights into emergent social behaviors and teamwork dynamics. Despite recent advances, there remains a significant gap in exploring how RL can be leveraged to further enhance the simulation of human behavior using LLMs, particularly in the context of web-based shopping environments.
@@
Reward design for RL. Reward design plays a central role in the effectiveness and generalization of RL algorithms, particularly in the context of aligning LLMs with desired behaviors. The prominent paradigm is Reinforcement Learning from Human Feedback (RLHF), which has been widely adopted to fine-tune LLMs using reward models trained on human preference data (Ouyang et al., 2022). While RLHF has demonstrated strong alignment capabilities, it is often bottlenecked by the high cost and limited scalability of collecting reliable human annotations (Touvron et al., 2023). Moreover, reward models themselves can introduce alignment biases and inaccuracies, especially when trained on limited or noisy preference comparisons (Gao et al., 2023). To alleviate these limitations, Direct Preference Optimization (DPO) (Rafailov et al., 2023) proposes a more efficient alternative that directly optimizes model parameters against human preference signals without an explicit reward model. Though computationally lighter, DPO and its variants still depend on the availability and quality of human-generated or approximated preference data, which can be inconsistent across tasks and domains. A complementary direction has emerged through Reinforcement Learning with Verifiable Rewards (RLVR), particularly suited for domains with deterministic correctness criteria such as code generation and mathematical reasoning (Guo et al., 2025; Su et al., 2025). RLVR frameworks employ rule-based verifiers to automatically compute reward signals based on strict correctness (e.g., exact string matching or functional equivalence) bypassing the need for human feedback. This shift toward automated objective reward functions has enabled the training of highly capable models such as DeepSeek-R1 (Guo et al., 2025) and inspired new policy optimization methods such as GRPO (Shao et al., 2024) and its recent extensions (Yu et al., 2025; Liu et al., 2025). Despite these advances, reward design remains a fundamental challenge in RL for human behaviors. RLHF offers flexibility for modeling subjective tasks, but often suffers from scalability and reliability issues (Alsagheer et al., 2025; Lee et al., 2023; Casper et al., 2023; Moskovitz et al., 2023). In contrast, RLVR provides high precision by relying on clearly defined evaluation criteria, but is limited to tasks where such criteria exist (Su et al., 2025; Mroueh, 2025; Wen et al., 2025; Lu et al., 2025c). To address the unique challenges of simulating human online shopping behavior, we propose a hybrid reward framework specifically tailored to this domain. For rationale generation, the framework leverages internal model signals to compensate for the absence of ground-truth rationales in natural settings. For action prediction, it adopts a hierarchical reward structure with difficulty-aware scaling, which mitigates reward hacking and alleviates the inefficiency caused by prolonged zero-reward feedback.
@@ 3 Methodology
In this section, we first formulate the problem of human behavior simulation in the context of web-based shopping. We then present the design of our proposed RL framework, Shop-R1, tailored specifically for simulating human behavior in this setting. Although user persona and explicit intention can influence human behavior, such information is neither collected nor observable in real-world e-commerce logs, nor can it be reliably reconstructed from HTML content or clickstream sequences. Crucially, the objective of our task is not to infer latent psychological states, but to model how real users behave given the same observable environment. Since human behavior is inherently non-deterministic, the goal is not to predict a single “correct” action; rather, it is to learn the distributional tendencies that characterize real user decision-making.
@@
Problem statement. In the context of web shopping, a user session is composed of a sequence of multi-step actions $a_{1\dots t\dots N}$, typically initiated with a search query and concluded by either a product purchase or a termination action (e.g., closing the browser). Following the setup of (Lu et al., 2025b), the action space comprises three primary action types: ‘type_and_submit’, ‘click’, and ‘terminate’. More details about the action space can be found in Appx. B. Each action $a_{t}$ is paired with a corresponding rationale $r_{t}$, which captures the user’s underlying motivation or rationale at that time step. The model also receives contextual information, i.e., the observation space, representing the current state of the web environment. This context is encoded as a simplified HTML structure, as introduced in (Lu et al., 2025e), which preserves essential layout and content elements while discarding non-informative components such as scripts and styles. The task of human online shopping behavior simulation is formally defined as learning a function $f$ that predicts the next rationale $r_{t}$ and action $a_{t}$, given the cumulative context $c_{1\dots t}$ and action history $a_{1\dots t-1}$:
$f(c_{1\dots t},a_{1\dots t-1})=r_{t},a_{t},$ (1)
@@
where $f$ denotes the model trained to simulate user behavior by generating the next-step rationale $r_{t}$ and action $a_{t}$ conditioned on the past context $c_{1\dots t-1}$, the corresponding past actions $a_{1\dots t-1}$, and the current context $c_{t}$. Thus, the model must jointly (a) understand the dynamic structure of the current web page, (b) integrate long-horizon action history, and (c) reason about the next plausible human action. This combination leads to long-horizon sequences and high-entropy free-form text predictions, making the task substantially more challenging than typical WebArena-style task-completion settings (Zhou et al., 2023). During the supervised fine-tuning stage, the rationales in training data are generated using LLMs and serve as supervision signals for a cold start. Need to note that no LLM-generated rationales are used during the subsequent RL stage.
@@
Cold start with SFT. Following the approach of (Guo et al., 2025), we initialize the behavior simulation model $f$ through supervised fine-tuning (SFT) on annotated trajectories, where each rationale is generated by Claude 3.5 Sonnet (Anthropic, 2024) via Amazon Bedrock, without leveraging any user profile information. This SFT phase acts as a cold start for subsequent RL, grounding the model in realistic rationale and action patterns. During this phase, the model is trained to jointly generate rationales and corresponding actions. The training objective is to maximize the likelihood of the ground truth rationale-action pairs, conditioned on the the input query $q_{t}=c_{1\dots t},a_{1\dots t-1},r_{1\dots t-1}$:
@@
Shop-R1. To better guide policy optimization in the human behavior simulation setting, we decompose each step into two sub-tasks: rationale generation and action prediction. Each sub-task is assigned a tailored reward to improve alignment and interpretability. To ensure the ease and correctness of parsing predicted rationales and actions from model outputs, we introduce a binary format reward, which encourages the model to produce responses in a structured JSON format. This format adheres to a dictionary schema with two keys: rationale and action. For rationale generation, we employ a self-certainty score (Kang et al., 2025; Zhao et al., 2025), which quantifies the model’s confidence in its generated rationale. Specifically, we compute the KL divergence between the model’s predictive distribution over the vocabulary and a uniform distribution, averaged over the entire output sequence:
@@
where $N$ is the number of tokens in the generated rationale $r_{t}$, $p_{ij}$ is the predicted probability of token $i$ at position $j$, and $U_{i}=\frac{1}{|V|}$ is the uniform distribution over the vocabulary $V$. Higher values of $s(\cdot)$ indicate greater certainty and consistency in the model’s reasoning.
@@
click 0.3 +0.2+0.2 (if name $\neq\varnothing$) $\text{DARS}\times\text{ROUGE-L}(\text{name})$
type_and_submit 0.3 +0.1+0.1 (if name $\neq\varnothing$) +0.1+0.1 (if text $\neq\varnothing$) $0.1\times\text{ROUGE-L}(\text{name}) +$$\text{DARS}\times\text{ROUGE-L}(\text{text})$
For action prediction, we replace the brittle binary signal with a hierarchical reward scheme that credits both the coarse‐grained action type and its fine‐grained sub-actions to stabilize training and discourage degenerate reward hacked policies. This hierarchical scheme densifies the reward landscape: it expands the set of profitable trajectories, lifts the agent out of the ‘no-reward’ plateau that typically stalls policy search, and makes reward hacking uneconomical. Concretely, every action, easy or hard, earns the same coarse-level reward once its high-level type is correct; only the more complex actions can unlock additional gains through their long-text sub-components. As a result, naively spamming the trivial ‘terminate’ action no longer yields a competitive payoff, while executing the full (‘click’, ‘type_and_submit’) sequence becomes the most lucrative strategy. Concretely, a ‘click’ action containing a sub-action, specifying the button name to be clicked; partial rewards are granted for the correctly predicted components. Likewise, ‘type_and_submit’ contains sub-action, providing the intended textual content. In contrast, ‘terminate’ has no sub-actions and is scored only at the action-type level. Prediction accuracy is measured with task-specific metrics: discrete action types use an exact‐match criterion, whereas free-form sub-actions are evaluated with ROUGE-L. A text-based sub-action, such as a button label or search query, earns a soft reward proportional to its ROUGE-L similarity to the ground truth, but only when that similarity exceeds a preset threshold (e.g., 0.75). Because long-text sub-actions are substantially harder, where modern webpages can expose thousands of candidate elements, we introduce a difficulty-aware reward scaling (DARS) factor that amplifies rewards for correctly predicting these components. This prevents reward hacking behaviors in which the agent repeatedly selects the trivial ‘terminate’ action to secure easy points. The proposed hierarchical reward scheme is summarized in Tab.˜1. Bringing these components together, the objective of Shop-R1is to maximize the combined reward signal derived from multiple sources, while regularizing with a KL divergence to a reference policy:
where $\pi_{\text{ref}}$ denotes a fixed reference policy, $v(a_{t})$ denotes the reward for action prediction and $\alpha$ and $\beta$ are hyperparameters that control the strength of the corresponding regularization terms. Need to note that because the action space used in our paper reflects general interaction patterns in Web GUIs, the proposed reward design can naturally generalize to different webpages and even to other simulation-oriented tasks involving Web GUI interactions. Importantly, the reward weights are not tuned for specific pages or environments. They function only as normalization constants to balance reward magnitudes, prevent trivial solutions (e.g., repeatedly predicting terminate), and ensure stable optimization rather than encoding task-specific biases.